Shelly is 3 years older than Michele. Four years ago the sum of their ages was 71. Find the age of each person now. (Example 3. Use the information in the table below showing algebraic expressions in terms of x.)

Answer by Maths68(1220)

Let
Present age of Shelly = x
Present age of Michele = x-3

Given
======
Four years ago
Shelly was = x-4
Michele was = (x-3)-4 = x-3-4 = x-7
Then
the sum of their ages was 71.
(x-4)+(x-7)=71
x-4+x-7=71
2x-11=71
2x=71+11
2x=82
2x/2=82/2
x=41

Present age of Shelly = x = 41 years old
Present age of Michele = x-3 = 41-3 = 38 years old



(Example 3. Use the information in the table below showing algebraic expressions in terms of x.)

No table given